OpenStudy (anonymous):

The Century club runs the concession stand during football games. They sell snacks and drinks. 50 snacks and 180 drinks will sell for $320. 75 snacks and 100 drinks will sell for $225. How much does each drink sell for?

OpenStudy (anonymous):

Let x represent the cost of a snack Let y represent the cost of a drink Our equations are gonna be in terms of cost since that is what we are interested in Now lets write a formula for both situations: 50x+180y=320 75x+100y=225 Now in order to solve for x and y you can use a multitude of methods. Which one you familiar with?/ Substitution, elimination, matrices?

OpenStudy (anonymous):

either substitution or elimination

OpenStudy (anonymous):

Lets do substitution... I think its easier and lets write the second equation in terms of y 75x+100y=225 100y=225-75x divide by a 100 \[y=\frac{225}{100}-\frac{75}{100}x\] y=2.25-.75x

OpenStudy (anonymous):

is it 1.5 ?

OpenStudy (anonymous):

Now replace the y in the first equation with 2.25-.75x So 50x+180y=320 Which becomes 50x+180(2.25-.75x)=320 Solve for x

OpenStudy (anonymous):

lol didnt have time to solve it .... Running out hmmmm @Jhannybean Can you finish this off for me I gotta run out and my family .... is gonna kill me

OpenStudy (anonymous):

Bump your quesion ... if not i will be back in 20 minutes

OpenStudy (anonymous):

i did.

OpenStudy (anonymous):

Thank you tho , ill be on .

OpenStudy (anonymous):

back

OpenStudy (anonymous):

50x-135x+405=320 -85x+405=320 -85x=-85 x=1 y=2.25-.75x y=2.25-.75(1) y=1.5 Yes you were correct